2x+5y=6000
3x+2y=6000
solve

Answer by TimothyLamb(4379)

2x + 5y = 6000
3x + 2y = 6000
---
5y = -2x + 6000
y = (-2/5)x + 1200
---
3x + 2y = 6000
2y = -3x + 6000
y = (-3/2)x + 3000
---
(-2/5)x + 1200 = (-3/2)x + 3000
(15/10)x + (-4/10)x = 1800
(11/10)x = 1800
x = (10/11)1800
---
y = (-2/5)x + 1200
y = (-2/5)(10/11)1800 + 1200
---
Answer:
You do the math:
x = (10/11)1800
y = (-2/5)(10/11)1800 + 1200
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php